Show that the curve with parametric equations


x = sin t and y = sin(t + sin t) for 0 ≤ t ≤ 2π


1
Expert's answer
2021-12-15T17:23:50-0500
"x_t'=\cos t, y'_t=\cos(t+\sin t)(1+\cos t)""y_x'=\dfrac{y_t'}{x_t'}=\dfrac{\cos(t+\sin t)(1+\cos t)}{\cos t}"

At the origin we have: "x=0" and "y=0."



"x=0=>\sin t=0=>t=\pi n, n\in \Z"

For "0\leq t \leq 2\pi" we have "t_1=0, t_2=\pi , t_3=2\pi."

Check for "y"



"t_1=0, y(0)=\sin (0+\sin(0)) =0, True""t_2=\pi, y(\pi)=\sin (\pi+\sin(\pi)) =0, True""t_3=2\pi, y(2\pi)=\sin (2\pi+\sin(2\pi)) =0, True"

So "y" is also satisfied. Put these values of "t" into "y_x'" to find the gradient of the tangent line:



"t_1=0,""y_x'|_{t=0}=\dfrac{\cos(0+\sin (0))(1+\cos (0))}{\cos (0)}=2""slope_1=m_1=2"

The equation of the tangent line is



"y-0=2(x-0)""y=2x""t_2=\pi,""y_x'|_{t=\pi}=\dfrac{\cos(\pi+\sin (\pi))(1+\cos (\pi))}{\cos (\pi)}=0""slope_2=m_2=0"

The equation of the tangent line is



"y-0=0(x-0)""y=0""t_3=2\pi,""y_x'|_{t=2\pi}=\dfrac{\cos(2\pi+\sin (2\pi))(1+\cos (2\pi))}{\cos (2\pi)}=0""slope_3=m_3=2=m_1"

The equation of the tangent line is



"y=2x"

Therefore our two tangents at the origin will be:



"y_1=2x, y_2=0"